Session: Feb 2000 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Morsiani at Fri Feb 11 17:59:37 MET 2000) Success

Shanghai (Huang Xinyong at Sat Feb 12 03:15:08 MET 2000) Success

Torun (K.Borkowski at Sun Feb 13 17:07:05 MET 2000) Success
No known problems.

Onsala (A. Polatidis at Mon Feb 14 10:15:30 MET 2000) Success
Lost 5 min around 0850 UT due to failure of polar motor (high winds). No other known problems.

Westerbork (Tony Foley at Mon Feb 14 10:46:31 MET 2000) Success
Used all 14 telescopes in tied array. No known problems

Effelsberg (A. Kraus at Tue Feb 15 08:37:08 MET 2000) Minor failures
Few scans had been lost due to insufficient slewing time and computer problems.

Jodrell (Alastair Gunn at Wed Feb 16 05:46:01 MET 2000) Minor failures
LO values not picked up by telescope at start of experiment. Correctly set manually at 06:23 UT. Late change in PIs schedule gave two tapes for this experiment. We did not record the last tape (which was not allocated) so the last half hour of data are missing.

Noto (C.Trigilio at Thu Feb 17 12:34:53 MET 2000) Minor failures
Wrong set of formatter till 8:12 UT. Data OK from 8:12 to end.

Calibration (Alastair Gunn at Wed Apr 19 12:57:29 METDST 2000) Success
Calibration data available from vlbeer: see file feb00/ga018.antab for details